If you are 30 m from the foot of the Leaning Tower of Pisa, the top of the Leaning Tower is at

65 ° angle of elevation.. If the tower were perpendicular to the ground, we would could form a

right right triangle and easily obtain the height.. Unfortunately, the tower is inclined towards us

4 ". Calculate to find the height of the tower with such data?


1
Expert's answer
2022-03-24T18:53:02-0400

Solution

To find the height of the tower we use sine rule which states

"\frac{a}{sineA}=\frac{b}{sineB}"


"\frac{30}{sine4°}=\frac{b}{sine111°}"


cross multiplication


"b=\frac{30sine111°}{sine4°}"


b=34.27m